OpenStudy (shadow004):

Medal and fan Which fractions are equivalent to ? Choose all answers that are correct. A. B. C. D.

OpenStudy (shadow004):

A. -10\25 B. -2\10 C. -4\20 D. 5\-25

OpenStudy (sooobored):

not enough information

OpenStudy (shadow004):

what else do you need

OpenStudy (danjs):

What is the given fraction in the question?

OpenStudy (shadow004):

-1\5

OpenStudy (sooobored):

are you familar with how to multiply fractions? and what happens when you multiply something by 1?

OpenStudy (shadow004):

no :'D

OpenStudy (shadow004):

i am not good with multiplication

OpenStudy (sooobored):

no to the first, no to the 2nd? no to both?

OpenStudy (shadow004):

what?

OpenStudy (sooobored):

i asked two questions

OpenStudy (shadow004):

i know

OpenStudy (shadow004):

plz help i have to get done by 12

OpenStudy (sooobored):

im going to assume you meant no to both my questions

OpenStudy (shadow004):

-_-#

OpenStudy (sooobored):

ok a fraction is made up of a numerator and a denominator the numerator is the top part of the fraction the denominator is the bottom part of the fraction \[\frac{x}{y}\] x would be the numerator y would be the denominator any questions?

OpenStudy (shadow004):

no...

OpenStudy (sooobored):

ok assuming you understand that, multiplication of fractions is very very easy, all you need to do is multiply the numerators together, and multiply the denominators together ex. \[\frac{a}{b}\times \frac{c}{d}= \frac{a\times c}{b\times d}\]

OpenStudy (shadow004):

k...

OpenStudy (sooobored):

now, what happens if you multiply any number by 1? well, 1x1 =1 2x1=2 3.92384 x1= 3.92384 12345678 x1 =12345678

OpenStudy (shadow004):

0_0.........................................

OpenStudy (sooobored):

multiply any number by 1 and you will get the same number

OpenStudy (shadow004):

the last 2

OpenStudy (shadow004):

i know that

OpenStudy (sooobored):

now, this also means that any number divided by itself is equal to 1

OpenStudy (shadow004):

i know that to

OpenStudy (sooobored):

ex. \[\frac{2}{2}=1\]\[\frac{100}{100}=1\] \[\frac{3.14159}{3.14159}=1\]

OpenStudy (shadow004):

i got that

OpenStudy (sooobored):

im not a mind reader, this is why i asked the questions i ahve no idea what you know

OpenStudy (sooobored):

now, we have -1/5 and we want to look for equivalent fractions

OpenStudy (shadow004):

-_- are you sure your not a mind reader

OpenStudy (sooobored):

if we multiply -1/5 by 1, then we should still get -1/5

OpenStudy (shadow004):

LOL

OpenStudy (sooobored):

that means if we multiply -1/5 by 2/2 (which is also 1) what do we get?

OpenStudy (shadow004):

but i got that

OpenStudy (shadow004):

and that

OpenStudy (sooobored):

what? i just went through all the theory above, now we're solving the problem

OpenStudy (shadow004):

just GOOOOOO i only have 30 more min

OpenStudy (sooobored):

we want to find equivalent fractions to -1/5 so by multiplying by 1 or say 2/2, we will still get a fraction that is equal to -1/5

OpenStudy (sooobored):

what is -1/5 times 2/2?

OpenStudy (sooobored):

\[\frac{-1}{5} \times \frac{2}{2}=?\]

OpenStudy (shadow004):

-3\3

OpenStudy (shadow004):

right?

OpenStudy (sooobored):

how is -1 x 2 equal to -3?

OpenStudy (sooobored):

and 2 x 5 equal to 3?

OpenStudy (shadow004):

-_- i told im not good

OpenStudy (sooobored):

as a suggestions, i suggest memorizing the "times table" or using a calculator but -1x2 = -2 2x5=10 so the fraction is?

OpenStudy (shadow004):

i did use a calculator

OpenStudy (sooobored):

then you're using it wrong since the only way to get 3 from 2 and 5 is by using the subtraction sign "-"

OpenStudy (shadow004):

i didn't tho

OpenStudy (sooobored):

the multiplication sign is "*" or "x"

OpenStudy (sooobored):

then you have a broken calculator

OpenStudy (shadow004):

I KNOW!!!!!!!!!!!!!!!

OpenStudy (sooobored):

use the one on your computer

OpenStudy (shadow004):

I JUST BROKE IT

OpenStudy (shadow004):

NOT THE COMPUTER

OpenStudy (sooobored):

i can teach fractions, i can teach the identity property, i cant teach multiplication....

OpenStudy (shadow004):

-_- i really dont have time please

OpenStudy (sooobored):

\[\frac{ -1}{5} \times \frac{2}{2} = \frac{-2}{10}\]

OpenStudy (sooobored):

-1/5 is equivalent to -2/10

OpenStudy (shadow004):

is that real answer

OpenStudy (sooobored):

thats one of the answer

OpenStudy (sooobored):

try multiplying -1/5 by 4/4 and by 5/5

OpenStudy (shadow004):

JUST HELP PLZ

OpenStudy (shadowlegendx):

Which fractions are equivalent to \[-\frac{ 1 }{ 5 }\] Reduce: A. -10\25 \[-\frac{ 10}{ 25 } \div \frac{ 5 }{ 5 } = -\frac{ 2 }{ 5 }\] B. -2\10 \[-\frac{ 2 }{ 10 } \div \frac{ 2 }{ 2 } = -\frac{ 1 }{ 5 }\] C. -4\20 \[-\frac{ 4 }{ 20 } \div \frac{ 4 }{ 4 } = - \frac{ 1 }{ 5 }\] D. 5\-25 \[-\frac{ 5 }{ 25 } \div \frac{ 5 }{ 5 } = - \frac{ 1 }{ 5 }\]

OpenStudy (shadow004):

I only have 9 min

OpenStudy (shadowlegendx):

Read, I showed you the answers, and how I got the answers.

OpenStudy (shadow004):

thank you so so much